suppose y varies directly as the square of x,
a. Find the constant of proportionality if y=4 when x=3. What is x when yis 16?
b. if the constant of variation is 5/8, what is y when x=6? what i x when y= 90?
c. find the constant of variation and the value of y when x=5 if y=2 when x=2

Answer by ewatrrr(24785)


Hi
y varies directly as the square of x
a. Find the constant of proportionality if y=4 when x=3. What is x when yis 16?
4 = k*9 k = 9/4
16 = 9/4*x*2
64/9 = x^2 x = ± 8/3
b. if the constant of variation is 5/8, what is y when x=6? what i x when y= 90?
y = (5/8)*36 = 22.5 90 = (5/8)x^2 x = ± 12
c. find the constant of variation and the value of y when x=5 if y=2 when x=2
2 = k*2^2 k = 2/4 = 1/2 y = (1/2)*25 y = 25/2 or 12.5